OpenStudy (catlover12345678):

Which proportion models this problem? On a soccer team, 5 out of 13 players surveyed said they had two or more siblings. The league has 520 players. Which is the best prediction of the number of players in the league that have two or more siblings please help

OpenStudy (debbieg):

5 out of 13 is a part (of the team) of out the whole (of the team) part(team)/whole(team) = part(league)/whole(league) You are given the total for the league; that's the "whole". The unknown is the "part" for the league. So set up the proportion, using a variable for the unknown.

OpenStudy (catlover12345678):

X over five hundred twenty equals five over thirteen am i right

OpenStudy (debbieg):

Yes, that's it!

OpenStudy (catlover12345678):

are you sure i just don't want to be wrong

OpenStudy (debbieg):

Yes, I'm sure.

OpenStudy (catlover12345678):

ok thanks

OpenStudy (ann1987):

There are 40 groups of 13 in the league. I found that by dividing 520 by 13. Each one of these 40 (13-member) groups has 5 people with 2 or more sibs. So there are about 200 people (40 × 5) with 2 or more siblings.

OpenStudy (catlover12345678):

The ratio of the perimeters of two rectangles is 5 to 8. The perimeter of the larger rectangle is 72 inches. What is the perimeter of the smaller rectangle?

OpenStudy (ann1987):

Set this up as a proportion: Small/Large = 5/8 = x/72 The product of the means (8 and x) is equal to the product of the extremes (5 and 72). So you can cross-multiply to solve this. 8x = 5 × 72 x = 5•72/8 x = 5×9 x = 45 Answer: The perimeter of the smaller rectangle is 45 inches.